three fourths of a certain number is 15 more than one fourth of the number.what are the numbers?

Answer by checkley79(3341)

3X/4=15+X/4
3X/4-X/4=15
2X/4=15
X/2=15
X=2*15
X=30 ANS.
PROOF:
3*30/4=15+30/4
90/4=15+30/4
90/4-30/4=15
60/4=15
15=15